You purchased stock nine years ago for $30,000 and sold it today for $83,190. What was the annual rate of return on your investment? Assume no dividends were paid.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

83190=30,000*(1+r/100)^9

(83190/30,000)^(1/9)=(1+r/100)

(1+r/100)=1.12

r=1.12-1

=12%